Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $115,700 of manufacturing overhead for an estimated allocation base of $89,000 direct material dollars to be used in production. The company has provided the following data for the just completed year: Purchase of raw materials $ 134,000 Direct labor cost $ 85,000 Manufacturing overhead costs: Indirect labor $ 127,500 Property taxes $ 8,600 Depreciation of equipment $ 16,000 Maintenance $ 11,000 Insurance $ 7,300 Rent, building $ 31,000 Beginning Ending Raw Materials $ 29,000 $ 11,000 Work in Process $ 49,000 $ 39,000 Finished Goods $ 72,000 $ 61,000 Required: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhead for the year. 3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials. 4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer. 5. Assume that the $39,000 ending balance in Work in Process includes $8,500 of direct materials. Given this assumption, supply the information missing below:
Answer #1

1) Overhead rate = 115700/89000 = 130% of direct material used

2) Applied overhead = (29000+134000-11000)*130% = $197600

Actual overhead = 127500+8600+16000+11000+7300+31000 = 201400

Under applied overhead = 197600-201400 = $3800

3) Schedule of cost of goods manufactured

Direct material used
Beginning direct material 29000
Add: Direct material purchase 134000
Less: Ending direct material -11000
Direct material used 152000
Direct labor 85000
Overhead applied 197600
Total manufacturing cost 434600
Beginning work in process 49000
Total Cost of work in process 483600
Less: Ending work in process -39000
Cost of goods manufactured 444600

4) Unadjusted cost of goods sold = 72000+444600-61000 = $455600

5) Ending work in process = 39000

Direct material = 8500

Applied overhead = (8500*1.3) = 11050

Direct labor = (39000-8500-11050) = $19450